OpenStudy (anonymous):

(x+2)^2/3 = 9 . HELLLLLP PLEASSE

OpenStudy (anonymous):

$$x = \sqrt{27} - 2$$

OpenStudy (anonymous):

how did you get that ?

OpenStudy (anonymous):

please explain

OpenStudy (anonymous):

(x+2)²/3 = 9 (x+2)² = 27 (x+2) = +/-sqrt(27) $$x = \pm \sqrt{27} - 2$$

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

the exponent is 2 over 3 just to let you know

OpenStudy (anonymous):

it is $$(x+2)^{2/3}$$ or $$((x+2)^2)/3$$ ?

OpenStudy (anonymous):

ohhh ok, sorry

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

let do again

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

(x+2)^(2/3) = 9 (x+2) = 9^(3/2) x = 27 -2 x = 25

OpenStudy (anonymous):

how did you get 27 from 9^(3/2) ?

OpenStudy (anonymous):

Ok, 9^(3/2) = $$\sqrt{9^3}$$ = $$\sqrt{729}$$ = 27

OpenStudy (anonymous):

Oh okay (: thank you . do you mind helping me with another ?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

(3x+9)^(3/4) = 81

OpenStudy (anonymous):

is it 27 ?

OpenStudy (anonymous):

or 6 i mean

OpenStudy (anonymous):

(3x+9)^(3/4) = 81 (3x+9) = 81^(4/3) x = (1/3) * (-9 + 243*3^(1/3)) x = -3+81*3^(1/3)

OpenStudy (anonymous):

wait so what is the answer?

OpenStudy (anonymous):

x = -3+81*3^(1/3)

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

$$x = -3 + 81\sqrt[3]{3}$$

OpenStudy (anonymous):

i dont understand how you got that

OpenStudy (anonymous):

ok, let's start again from here: (3x+9) = 81^(4/3) $$3x+9 = 243\sqrt[3]{3}$$ $$3x = -9 + 243\sqrt[3]{3}$$ $$x = -3 + 81\sqrt[3]{3}$$

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhhhhh

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

wait how did you get 243?

OpenStudy (anonymous):

81^(4/3) = $$(\sqrt[3]{81})^4 = (3\sqrt[3]{3})^4 = 243\sqrt[3]{3}$$

OpenStudy (anonymous):

oh okay i understand now

OpenStudy (anonymous):

or you can see in this way: $$\sqrt[3]{81^4} = 243\sqrt[3]{3}$$

OpenStudy (anonymous):

got it?